OpenStudy (hatadtadatd):

An acute angle θ is in a right triangle with sin θ = five sixths. What is the value of cot θ?

OpenStudy (hatadtadatd):

@agent0smith

OpenStudy (hatadtadatd):

@sammixboo

OpenStudy (hatadtadatd):

@jigglypuff314

OpenStudy (hatadtadatd):

@mayankdevnani

OpenStudy (aakashtomar):

Hmm. First take out cos of theta by the eqn cosx= (√1-sin^2x) Then use cosx/sinx = cotx

OpenStudy (hatadtadatd):

I've tried that but I always mess it up

OpenStudy (hatadtadatd):

I kept getting 5/6 but that's not an answer Answer: five divided by the square root of eleven six divided by the square root of eleven square root of eleven divided by five square root of eleven divided by six

OpenStudy (agent0smith):

Then try making a right triangle. Remember sin theta = opposite/hypotenuse you have sin theta = 5/6

OpenStudy (hatadtadatd):

..I'm confused I've tried

OpenStudy (agent0smith):

sin theta = opposite/hypotenuse you have sin theta = 5/6 So draw a right triangle with an angle theta. Tell me what the opposite is, and what the hypotenuse is, based on what I wrote above.

OpenStudy (hatadtadatd):

I don't know this, you can go haha

OpenStudy (agent0smith):

I'll make it even easier sin theta = opposite/hypotenuse sin theta = 5 / 6 Tell me what the opposite is, and what the hypotenuse is.

OpenStudy (hatadtadatd):

opposite is 5 and hypotenuse is 6

OpenStudy (agent0smith):

So make a triangle like this http://www.technologyuk.net/mathematics/trigonometry/images/trigonometry_0018.gif

OpenStudy (hatadtadatd):

Alright

OpenStudy (hatadtadatd):

I did that

OpenStudy (phi):

the next step is use pythagoras to find the 3rd side.

OpenStudy (hatadtadatd):

I'm so confused with this whole thing

OpenStudy (phi):

do you know how to do that ? \[ a^2 + b^2 = c^2 \] where a and b are the "legs" and c is the hypotenuse

OpenStudy (agent0smith):

 Find b

OpenStudy (hatadtadatd):

Alright, I did that

OpenStudy (hatadtadatd):

Is it A?

OpenStudy (agent0smith):

Show work. I'll check your work. I won't check whether it's A or not.

OpenStudy (hatadtadatd):

@phi

OpenStudy (agent0smith):

Don't tag someone else to get them to check your answer (which could very easily be a guess) Show work and prove to us that you did it correct.